OpenStudy (anonymous):

What happens to the graph of y=f(x) when x is replaced with x/3, -x, 3x

OpenStudy (anonymous):

\[y=f(-x)\] looks like \(y=f(x)\) reflected across the \(y\) axis

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\(y=f(\frac{x}{3})\) looks like \(y=f(x)\) only stretched out by a factor of 3 as x goes from say 0 to 9, x/3 goes only from 0 to 3

OpenStudy (anonymous):

reverse for \(y=f(3x)\) try graphing \(y=x^3-x\) and \(y=(3x)^3-3x\) and compare http://www.wolframalpha.com/input/?i=%28x%2F3%29^3%2Bx%2F3%2C+x^3-x%2C+%283x%29^3-3x+domain+-2..2

OpenStudy (anonymous):

alright awesome, thank you!!